10) When 10.2 g of canola oil at 25.0°C is placed in a wok, 3.34 kJ of heat is required to heat it to a temperature of 196.4°C. What is the specific heat of canola oil?



1
Expert's answer
2021-10-29T02:18:31-0400

c=Q/(m"\Delta"T)

mass=(10.2/1000)kg

"\Delta" T=(196. 4-25.0)=171.4K

c=3.34×1000 ×1000 /(171.4×10.2)

c=1910.4J/(kgK)